OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = (x – 5)^2 + 10?

sam (.sam.):

Range = 10

sam (.sam.):

Domain is all real numbers Range is 10

Directrix (directrix):

Domain: the set of Real Numbers Range: y ≥ 10

OpenStudy (anonymous):

Domain: \[\mathbb{R}=(-\infty, \infty)\] Range \[[10,\infty)\]